Prove that tan inverse 1+tan inverse2 +tan inverse3=π

Grade:12

1 Answers

amit kumar
6 years ago
If u know that Tan inverse A +Tan inverse B=π+Tan inv(A+B)/1-AB if AB>1 and A,B>0Then simplyTan(inv)2+Tan(inv)3=π+Tan(inv)(2+3)/(1-2.3)=π+Tan(inv)-1=π-Tan(inv)1So tan(inv)1+tan(inv)2+tan(inv)3=π